A regression study of the relationship between the boiling temperature of water (y,in degrees Fahrenheit) and the atmospheric pressure (x, in inches of mercury) based on a data set withn= 30 observations resulted in the following regression equation:ˆy= 149 + 2.1x. If the average of the 30 atmospheric pressure reading (i.e. ̄x)was 20 inches of mercury, then what the average of the boiling temperature reading(i.e. ̄y) for these 30 observations.
Answer #1

The coefficient b0 = (Sum y - b1 * Sum x) / n

Here b0 = 149, b1 = slope = 2.1

Sum x = Mean * Total observations = 20 * 30 = 600

n = 30

Therefore (Sum y - 2.1 * 600) / 30 = 149

Solving, we get : Sum y = (149 * 30) + (2.1 * 600) = 4470 + 1260

Sum y = 5730

Average y = = 5730 / 30 = 191